# Sample Compression for Self-Certified Continual Learning

### v3

@@ 2 Background
$\mathcal{L}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu))\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,{}_{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}}\quantity(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)),\frac{1}{n-\operatorname{|\mathbf{i}|}}\quantity[\log\!\matrixquantity(n\\ \operatorname{|\mathbf{i}|})\!\!+\!\log\!\quantity(\frac{2\sqrt{n-\operatorname{|\mathbf{i}|}}}{\zeta(\operatorname{|\mathbf{i}|})P_{\mathscr{M}(\operatorname{\mathbf{i}})}(\mu)\delta})])\,,$
with $\mathrm{kl}(q,p)=q\ln\frac{q}{p}+(1-q)\ln\frac{1-q}{1-p}$ the binary Kullback-Leibler divergence and its inverse
$\mathrm{kl}^{-1}(q,\varepsilon)=\mbox{\rm argsup}_{p\in[0,1]}\left\{\mathrm{kl}(q,p)\leq\varepsilon\right\}.$ (5)
On the one hand, given a fixed compression set size $\operatorname{|\mathbf{i}|}$, the bound decreases when the training set size $n$ increase. On the other hand, given a fixed $n$, the bound increases with $\operatorname{|\mathbf{i}|}$.
Pick-to-Learn.  To obtain sample compression guarantees for a class of predictors, it is necessary to prove that the learned predictor only depends on a small subset of the data and (optionally) a message. Some predictors, such as the SVM boser_svm and the perceptron rosenblatt1958perceptron ; moran2020perceptron , have straightforward compression scheme, as they only depend on a subset of the data after training. Some algorithms, such as the set covering machine (SCM) marchand2002set ; marchand2003set ; marchand2005learning and decision trees shah_sample_2007 , necessitate hand-crafted compression schemes involving a message. Up until recently, there was no sample compression scheme for deep neural networks.
Algorithm 1 Original Pick-To-Learn (P2L) paccagnan_pick_learn_2023
@@
0:  $\theta_{\rm init}$Initial parameters of the model $f_{\theta}$}
@@
The meta-algorithm Pick-To-Learn (P2L) was proposed by Paccagnan and Campi paccagnan_pick_learn_2023 as a compression scheme for any class of predictors, with a specific focus to deep neural networks. The meta-algorithm acts as an aditional training loop, by iteratively selecting datapoints over which the model is updated. Starting with initial parameters $\theta_{\rm init}$, P2L evaluates the predictor on the whole training dataset, adds the datapoints with the highest losses to the compression set and then updates the predictor using the compression set. The meta-algorithm stops once the losses of training examples not in the compression set are lower than a given stopping criteria ($-\ln(0.5)$ for the cross-entropy loss, which is equivalent to achieving zero errors on the complement set $S^{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}$). We provide the pseudo-code of the original P2L in Algorithm 1. Recent works introduce improvements to the original scheme, notably early stopping strategies marks2025pick2learn-gp ; bazinet2024 and extension to the Bayesian setting marks2025pick2learn-gp .
@@ 3 Sample Compression for Continual Learning
A striking realization coming from the P2L algorithm is that only a small fraction of the training set – the compression set – needs to be provided to the learner in order to achieve good generalization. The winning strategy is to select this compression set to ensure a low risk on the training samples not being retained in the compression set (i.e., the complement of the compression set). This observation motivates our strategy for managing the replay buffer of our new algorithm Continual Pick-To-Learn (CoP2L): we subsample the complement set to create the replay buffer, in order to maintain a low risk on these examples while learning subsequent tasks. When needed, CoP2L can choose to add a datapoint from the buffer to the compression set, thus mitigating the forgetting by focusing on a well-chosen datapoint.
@@
13:  return $\theta_{k},C_{k}$. {Final parameters and compression set}
@@
Each new task is learned by a modified version of the P2L meta-algorithm (Algorithm 2, entitled mP2L), as explained further down. That is, for each new task, the proposed continual learning algorithm (Algorithm 3, entitled CoP2L) calls mP2L and updates the buffer using the datapoints that were not chosen in the compression set. Similarly to the replay buffer method, at the end of each task $t$, the buffer contains $\lfloor\tfrac{m}{t}\rfloor$ datapoints of each previous task, with $m$ the maximum buffer size.
We modified Pick-To-Learn in two significant ways. (i) First, we introduced weights to the loss functions to tackle the imbalance problem between the current task and the previous tasks. When working with a replay buffer, the class imbalance problem is taken care of by training on a number of datapoints from previous tasks. As our model is only trained on a very limited subset of the data, we need to find another way to mitigate the effect of class imbalance. Before starting the training on a new task, we set the weight of the datapoints from the previous task to $\omega$ and the weight of the current task to $1$. For the stopping criteria of mP2L to be satisfied, the worst loss on the current task must be less than $\gamma$ and the worst loss on the previous task must be less than $\tfrac{\gamma}{\omega}$. This ensures that the model has achieved perfect accuracy on the current task and all the datapoints from the previous tasks that are found in the buffer. (ii) The second modification is the choice of the returned predictor by mP2L. The original P2L algorithm trains the model until it first achieves zero errors on the complement set. In presence of a noisy dataset, this might lead to an overfitting behavior, P2L adding most training samples into its compression set, which in turn gives a large generalization bound value. Instead, we perform early stopping based on the bound value (as proposed by bazinet2024 ). That is, mP2L relies on the trade-off encoded in Theorem 2.1 between the accuracy on the complement set and the complexity of the model. More precisely, it returns the model’s parameters $\theta$ that minimizes
$\Psi(S,\theta,C)$ $=\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S\setminus C}\quantity(\theta),\ \tfrac{1}{|S\setminus C|}\log\quantity(\tfrac{2\sqrt{|S\setminus C|}{|S|\choose|C|}}{\zeta(|C|)\delta})),$ (6)
with $\mathrm{kl}^{-1}$ defined by Equation 5. Thanks to Theorem 2.1, we have that $\operatorname{\mathcal{L}}_{\mathcal{D}}(\theta)\leq\Psi(S,\theta,C)$ with probability at least $1-\delta$.
Note that for the first observed task, the mP2L procedure starts from randomly initialized parameters $\theta_{0}$ and is executed on the training sample $S_{1}$ (as it is done by the original P2L). Then, for every subsequent task $t\in\{2,\ldots,T\}$, the mP2L procedure is initialized to the previously learned parameters $\theta_{t-1}$. It then learns from the current task sample $S_{t}$ and a subset of randomly selected instances from previous tasks. The latter is obtained from the procedure sample$(S,m)$ (see Algorithm 3), which represents the random sampling of $m$ instances of $S$ without replacement.
@@
Compression and reconstruction scheme.  From a theoretical standpoint, in contrast to the original P2L, both mP2L and CoP2L cannot be reconstructed straightforwardly without a message for two reasons: (1) when only the compression set is given as input, the bound function $\Psi(S,\theta,C)$ used as stopping criteria in mP2L cannot be computed on the whole dataset; (2) the sampling procedure sample$(S,m)$ in CoP2L cannot be reproduced in the reconstruction step. In Appendix A, we provide the compression and reconstruction functions of CoP2L, which give rise to sample compression bounds as presented in Theorem 3.1. Noteworthy, after learning on $T$ tasks, the compression function provides two sample compression sets $S^{\operatorname{\mathbf{i}}}$ and $S^{\operatorname{\mathbf{j}}}$, along with a message pair $(\mu_{1},\mu_{2})$ chosen among the set of all possible messages, denoted as $\mathscr{M}(\operatorname{\mathbf{j}},T)=\{2,\ldots,T\}^{|\operatorname{\mathbf{j}}|}\times\left[\boldsymbol{\times}_{t=1}^{T}\{1,\ldots,n_{t}+|{\mathcal{B}}|\}\right].$ As explained further in appendix, the use of two compression sets, along with the message $\mu_{1}=[\mu_{1}^{i}]_{i=1}^{T}$, serves to the proper treatement of the buffer $B^{\star}$, while the message $\mu_{2}=[\mu_{2}^{i}]_{i=1}^{T}$ contains the number of iterations $K^{\star}$ to perform for each call to mP2L.
@@
Generalization bound.  Given the compression and reconstruction function detailed in Apprendix A, Theorem 2.1 can be used to obtain generalization bounds on the last learned task, which is done by mP2L to compute the bound $\Psi$. However, this bound holds for only one distribution of data. Therefore, it cannot be applied to bound the risks of the learned predictor on previously tasks seen by CoP2L. The next Theorem 3.1 holds for any previously learned tasks. For a dataset $S_{t}\sim\operatorname{\mathcal{D}}_{t}$, we denote the loss on the complement set of the joint set $(S_{t}^{\operatorname{\mathbf{i}}}\cup S_{t}^{\operatorname{\mathbf{j}}})^{\rm\bf c}$ as ^$\widehat{{\mathcal{L}}}_{S_{t}}^{\ \operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\!(\theta)$. Moreover, we denote the reconstruction function of CoP2L as $\operatorname{\mathscr{R}}_{1:T}(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)=\operatorname{\mathscr{R}}\big{(}S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu\mid S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}\big{)}.$ This formulation is important for the following theorem, as the reconstruction function for task $t$ is conditioned on all datasets $S_{1}$ to $S_{T}$, with the exception of $S_{t}$.
@@
For any set of independent distributions $\left\{{\mathcal{D}}_{t}\right\}_{t=1}^{T}$ over ${\mathcal{X}}\times{\mathcal{Y}}$, with $\operatorname{\mathscr{R}}_{1:T}$ the reconstruction function of CoP2L (Algorithm 5), with the family of set of messages $\{\mathscr{M}(\operatorname{\mathbf{j}},T)|\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)\}$ defined for CoP2L, for any loss $\ell:\Theta\times{\mathcal{X}}\times{\mathcal{Y}}\to[0,1]$, and for any $\delta\in(0,1]$, with probability at least $1-\delta$ over the draw of $S_{t}\sim{\mathcal{D}}_{t},t=1,\ldots,T$, we have
$\forall t\in\{1,\ldots,T\},{\mathbf{i}}\in\operatorname{\mathscr{P}}(n_{t}),\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n_{t}),\mu=(\mu_{1},[\mu_{2}^{i}]_{i=1}^{T})\in\mathscr{M}(\operatorname{\mathbf{j}},T):$
$\mathrm{kl}\quantity(\widehat{{\mathcal{L}}}_{S_{t}}^{\ \operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\!\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)),\ {\mathcal{L}}_{\operatorname{\mathcal{D}}_{t}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)))$
$\leq\tfrac{1}{n_{t}-|{\mathbf{i}}|-|\mathbf{j}|}\big{[}\log{\textstyle{n_{t}\choose|{\mathbf{i}}|}}+\log{\textstyle{n_{t}-\operatorname{|\mathbf{i}|}\choose|\mathbf{j}|}}+\textstyle\sum_{i=1}^{T}\log\tfrac{1}{\zeta(\mu_{2}^{i})}+|\mathbf{j}|\log(T-1)+\log\tfrac{T}{\zeta(|{\mathbf{i}}|)\zeta(|\operatorname{\mathbf{j}}|)\delta}\big{]}\,.$
The proof is given in Appendix C. It first relies on a tighter version of Theorem 2.1, where the $2\sqrt{n-\operatorname{|\mathbf{i}|}}$ term is removed and a second compression set is added. Then, this new result is adapted to the continual learning setting. For all tasks $1$ through $T$, this new theorem encodes a tradeoff similar to Theorem 2.1 between the predictor accuracy and its complexity. Here, the predictor complexity is evaluated using the size of the compression sets $\operatorname{\mathbf{i}}$ and $\operatorname{\mathbf{j}}$ and the probability of choosing a message. The probability of the messages $\mu_{1}$ and $\mu_{2}$ are functions of the size of $\operatorname{\mathbf{j}}$ and the number of tasks. Indeed, the probability of $\mu_{1}$ decays when $|\mathbf{j}|$ and $T$ grow larger. Moreover, the probability of $\mu_{2}$ depends on the size of the compression set outputted by mP2L at each task $T$. Thus, the larger the compression sets, the smaller the probability of both messages. In conclusion, the bound encodes the ability of a model to minimizes the loss on the dataset whilst still keeping $\operatorname{\mathbf{i}}$ and $\operatorname{\mathbf{j}}$ the small.

### v4

@@ 2 BACKGROUND
$\mathcal{L}_{\operatorname{\mathcal{D}}}\quantity(\theta_{\operatorname{\mathbf{i}},\mu})\leq\mathrm{kl}^{-1}\left(\operatorname{\widehat{\mathcal{L}}}_{S}^{{\;\;}_{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}}\quantity(\theta_{\operatorname{\mathbf{i}},\mu}),\frac{1}{\operatorname{|\mathbf{i}^{{\rm\bf c}}|}}\ln\frac{2\smallmatrixquantity(n\\ \operatorname{|\mathbf{i}|})\sqrt{\operatorname{|\mathbf{i}^{{\rm\bf c}}|}}}{\zeta(\operatorname{|\mathbf{i}|})P_{\mathscr{M}(\operatorname{\mathbf{i}})}(\mu)\delta}\right),$
with $\theta_{\operatorname{\mathbf{i}},\mu}\,{=}\,\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)$ the reconstructed parameters, $\mathrm{kl}(q,p)=q\ln\frac{q}{p}+(1-q)\ln\frac{1-q}{1-p}$ the binary Kullback-Leibler divergence and its inverse
$\mathrm{kl}^{-1}(q,\varepsilon)=\operatorname*{arg\,sup}_{p\in[0,1]}\left\{\mathrm{kl}(q,p)\leq\varepsilon\right\}.$ (4)
On the one hand, given a fixed compression set size $\operatorname{|\mathbf{i}|}$, the bound decreases when the training set size $n$ increases. On the other hand, given a fixed dataset size $n$, the bound increases with $\operatorname{|\mathbf{i}|}$.
Pick-to-Learn.
To obtain sample compression guarantees for a class of predictors, it is necessary to prove that the learned predictor only depends on a small subset of the data and (optionally) a message. Some predictors, such as the SVM (Boser et al., 1992) and the perceptron (Rosenblatt, 1958; Moran et al., 2020), have a straightforward compression scheme, as they only depend on a subset of the data after training. Some learning algorithms, such as the set covering machine (SCM) (Marchand and Shawe-Taylor, 2002; Marchand et al., 2003; Marchand and Sokolova, 2005) and decision trees (Shah, 2007), necessitate hand-crafted compression schemes involving a message. Up until recently, there was no sample compression scheme for deep neural networks.
Algorithm 1 Pick-To-Learn (P2L)    (Paccagnan et al., 2024)
0:  $\theta_{\rm init}$ {Initial parameters of the model $f_{\theta}$}
@@
The meta-algorithm Pick-To-Learn (P2L) was proposed by Paccagnan et al. (2024) as a compression scheme for any class of predictors, with a specific focus on deep neural networks. The meta-algorithm acts as an additional training loop, by iteratively selecting datapoints over which the model is updated. Starting with initial parameters $\theta_{\rm init}$, P2L evaluates the predictor on the whole training dataset, adds the datapoints with the highest losses to the compression set and then updates the predictor using the compression set. The meta-algorithm stops once the losses of training points that do not belong to the compression set are lower than a given stopping criterion $\gamma$. We provide the pseudo-code of P2L in Algorithm 1. Recent works introduce improvements to the original scheme, notably early stopping strategies (Bazinet et al., 2025; Paccagnan et al., 2025) and extension to the Bayesian setting (Marks and Paccagnan, 2025).
@@ 3 SAMPLE COMPRESSION FOR CONTINUAL LEARNING
A striking realization coming from the P2L algorithm is that only a small fraction of the training set – the compression set – needs to be provided to the learner in order to achieve good generalization. The winning strategy is to select this compression set to ensure a low risk on the training samples not being retained in the compression set (i.e., the complement of the compression set). This observation motivates the strategy employed in our new algorithm Continual Pick-To-Learn (CoP2L), which manages the replay buffer by subsampling only from the complement set instead of the whole training set. Then, while learning subsequent tasks, CoP2L maintains a low risk on previous tasks by adding well-chosen datapoints from the buffer to its compression set. This strategy inherently both mitigates the forgetting and enables the computation of sample compression bounds.
@@
13:  return $\theta_{k},C_{k}$ {Final parameters and compression set}
In CoP2L, each new task is learned by a modified version of the P2L meta-algorithm (Algorithm 2, entitled mP2L), as explained further down. That is, for each new task, the proposed continual learning algorithm (Algorithm 3) calls mP2L and then updates the replay buffer by sampling datapoints that were not chosen in the compression set. Similar to the replay buffer method, at the end of each task $t$, the buffer contains $\lfloor\tfrac{m}{t}\rfloor$ datapoints of each previous task, with $m$ the maximum buffer size.
We modified Pick-To-Learn in two significant ways. (i) First, we introduced weights to the loss functions to tackle the imbalance problem between the current task and the previous tasks. Traditionally, when working with a replay buffer, the class imbalance problem is taken care of by training on a number of datapoints from previous tasks. As our model is only trained on a very limited subset of the data, another strategy must be implemented to mitigate the class imbalance effect. Before starting the training on a new task, mP2L sets the weight of the datapoints from the previous task to $\omega\,{>}\,1$ and the weight of the current task to $1$. For the stopping criteria of mP2L to be satisfied, the worst loss on the current task must be less than $\gamma$ and the worst loss on the previous tasks must be less than $\tfrac{\gamma}{\omega}$. (ii) The original P2L algorithm trains the model until it first achieves zero errors on the complement set. Instead, mP2L performs early stopping based on the bound value (as proposed by Bazinet et al., 2025). That is, mP2L relies on the trade-off encoded in Theorem˜2.1 between the accuracy on the complement set $C$ and the complexity of the model $f_{\theta}$. More precisely, it returns the model’s parameters $\theta$ that minimizes
$\Psi(S,\theta,C)\!=\!\mathrm{kl}^{-1}\!\left(\!\operatorname{\widehat{\mathcal{L}}}_{S\setminus C}(\theta),\tfrac{1}{|S\setminus C|}\ln\tfrac{2\sqrt{|S\setminus C|}{|S|\choose|C|}}{\zeta(|C|)\delta}\right)\!,\!$ (5)
with $\mathrm{kl}^{-1}$ given by Eq. (4). Thanks to Theorem˜2.1, we have that $\operatorname{\mathcal{L}}_{\mathcal{D}}(\theta)\leq\Psi(S,\theta,C)$ with high probability.
Note that for the first observed task, the mP2L procedure starts from randomly initialized parameters $\theta_{0}$ and is executed on the training sample $S_{1}$ (as it is done by the original P2L). Then, for every subsequent task $t\in\{2,\ldots,T\}$, the mP2L procedure is initialized to the previously learned parameters $\theta_{t-1}$. It then learns from the current task sample $S_{t}$ and a subset of randomly selected instances from previous tasks. The latter is obtained from the procedure sample(S,m)(S,m) (see Algorithm 3), which represents the random sampling of $m$ instances of $S$ without replacement. [^2] Although we only discuss random sampling, any buffer management technique applied to the complement set would be valid in CoP2L, such as coreset methods.
@@
Compression and reconstruction scheme.
@@
From a theoretical standpoint, in contrast to the original P2L, both mP2L and CoP2L cannot be reconstructed straightforwardly without a message for two reasons: (1) when only the compression set is given as input, the bound function $\Psi(S,\theta,C)$ used as stopping criteria in mP2L cannot be computed on the whole dataset; (2) the sampling procedure sample(S,m)(S,m) in CoP2L cannot be reproduced in the reconstruction step. In Appendix B, we prove that CoP2L is a sample-compression algorithm by providing its compression and reconstruction functions, which subsequently gives rise to sample-compression bounds as presented in Theorem˜3.1. Noteworthy, after learning on $T$ tasks, the compression function provides two compression sets $S^{\operatorname{\mathbf{i}}}$ and $S^{\operatorname{\mathbf{j}}}$, along with a message pair $\mu=(\mu_{1},\mu_{2})$. As explained further in the appendix, the use of two compression sets, along with the first part of the message $\mu_{1}=[\mu_{1}^{i}]_{i=1}^{T}$, serves the proper treatment of the weighted buffer $B^{\star}$, while the second part of the message $\mu_{2}=[\mu_{2}^{i}]_{i=1}^{T}$ contains the number of iterations $K^{\star}$ to perform for each call to mP2L. The set of all possible message pairs is denoted as $\mathscr{M}_{1:T}(\operatorname{\mathbf{j}})=\{2,\ldots,T\}^{|\operatorname{\mathbf{j}}|}\times\left[\boldsymbol{\times}_{t=1}^{T}\{1,\ldots,n_{t}+|{\mathcal{B}}|\}\right].$
Generalization bound.
@@
Given the compression and reconstruction function detailed in Appendix B, Theorem˜2.1 can be used to obtain generalization bounds on the last learned task, which is done by mP2L to compute the bound $\Psi$. However, this bound holds for only one distribution of data. Therefore, it cannot be applied to bound the risks of the learned predictor on previous tasks seen by CoP2L.
The next Theorem 3.1 holds for any previously learned tasks. For a dataset $S_{t}\sim\operatorname{\mathcal{D}}_{t}$, we denote the loss on the complement set of the joint set $S_{t}^{\operatorname{\mathbf{i}}}\cup S_{t}^{\operatorname{\mathbf{j}}}$ as $\widehat{{\mathcal{L}}}_{S_{t}}^{\ \operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\!(\theta)$. Moreover, we denote the reconstruction function of CoP2L as $\operatorname{\mathscr{R}}_{1:T}\big(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu\mid S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}\big).$ This formulation is important for the following theorem, as the reconstruction function for task $t$ is conditioned on all datasets $S_{1}$ to $S_{T}$, with the exception of $S_{t}$.
@@
For any set of distributions $\left\{{\mathcal{D}}_{t}\right\}_{t=1}^{T}$ over ${\mathcal{X}}\times{\mathcal{Y}}$ sampled from $\operatorname{\mathfrak{D}}$, and for any $\delta\in(0,1]$, with probability at least $1-\delta$ over the draw of $S_{t^{\prime}}\sim{\mathcal{D}}_{t^{\prime}},{t^{\prime}}=1,\ldots,T$, we have
$\forall t\in\{1,\ldots,T\},\ {\mathbf{i}},\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n_{t}),\ \mu\!=\!(\mu_{1},\mu_{2})\in\mathscr{M}_{1:T}(\operatorname{\mathbf{j}}):$
${\mathcal{L}}_{\operatorname{\mathcal{D}}_{t}}\quantity(\theta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}^{(t)})\leq\mathrm{kl}^{-1}\left(\widehat{{\mathcal{L}}}_{S_{t}}^{\ \operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\!\quantity(\theta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}^{(t)}),\frac{\epsilon(\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu)}{n_{t}-|{\mathbf{i}}|-|\mathbf{j}|}\right),$
with $\theta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}^{(t)}=\operatorname{\mathscr{R}}_{1:T}\big(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu\,{\mid}\,S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}\big)$ the reconstructed parameters given by Algorithm˜5 and
$\epsilon(\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu)=\ln\left[\tfrac{T}{\delta}\textstyle{n_{t}\choose\operatorname{|\mathbf{i}|}}\textstyle{n_{t}-\operatorname{|\mathbf{i}|}\choose|\mathbf{j}|}\tfrac{(T-1)^{|\mathbf{j}|}}{\zeta(|{\mathbf{i}}|)\zeta(|\operatorname{\mathbf{j}}|)}\prod_{i=1}^{T}\tfrac{1}{\zeta(\mu_{2}^{i})}\right]\!.$
The proof is given in Appendix˜D. It first relies on a tighter version of Theorem˜2.1, where the $2\sqrt{n-\operatorname{|\mathbf{i}|}}$ term is removed and a second compression set is considered. This new result is tailored to the continual learning setting. For all tasks $1$ through $T$, Theorem˜3.1 encodes a tradeoff similar to Theorem˜2.1 between the predictor accuracy and its complexity. Here, the predictor complexity is evaluated using the size of the compression sets $\operatorname{\mathbf{i}}$ and $\operatorname{\mathbf{j}}$ and the probability of choosing a message. The probability of the messages $\mu_{1}$ and $\mu_{2}$ are functions of the size of $\operatorname{\mathbf{j}}$ and the number of tasks. Indeed, the probability of $\mu_{1}$ decays when $|\mathbf{j}|$ and $T$ grow larger. Moreover, the probability of $\mu_{2}$ depends on the size of the compression set outputted by mP2L at each task $T$. Thus, the larger the compression sets, the smaller the probability of both messages. In conclusion, the bound encodes the ability of a model to minimize the loss on the dataset whilst still keeping $\operatorname{\mathbf{i}}$ and $\operatorname{\mathbf{j}}$ small.